User name Carlos
Log entry time 16:21:03 on October16,2004
Entry number 134429
keyword=Trigger rates falling down: reset procedure
Trigger rates went down from 50Hz to 25Hz after run 3789.
Two possible causes:
1. You should go and see the cpuvme window (snapshot below) of ROC17 in adaql2 (usually in the desktop next to CODA runcontrol). You can type Data(). If the ouput does not look as the one below (last three rows of PMT table equal to 0), you should power-cycle the trigger (procedure in post-it next to adaql2).
2. If everything looks ok, pedestals might be wrong. To correct that, stop the run and type in the ROC17 window:
The next run should have the usual rates now.
The whole procedure took between 15:30 and 16:03. Should be quicker next time.